OpenStudy (anonymous):

evaluate integral (sin^2(t) dt) from 0 to 2pi?

OpenStudy (science0229):

\[\int\limits_{0}^{2\pi}\sin^2(t)dt\]

OpenStudy (amistre64):

sin^2 has a simpler version ... consider the identity of cos(2t) to find a suitable substution for sin^2(t)

OpenStudy (science0229):

\[\cos(2t)=1-2\sin^2(t)\rightarrow \sin^2(t)=\frac{ 1-\cos(2t) }{ 2 }\]

OpenStudy (science0229):

That is called power reduction formula. It's good to know them to make integration easier.

OpenStudy (science0229):

Can you take it from here?

OpenStudy (anonymous):

ys

OpenStudy (science0229):

good luck

OpenStudy (science0229):

:)

OpenStudy (anonymous):

\[\int\limits \sin ^2(t) \, dt=\frac{t}{2}-\frac{1}{4} \sin (2 t) \]\[\int\limits_0^{2 \pi } \sin ^2(t) \, dt=\pi \]\[\left(\frac{t}{2}-\frac{1}{4} \sin (2 t)\text{/.}\, t\to 2 \pi \right)-\left(\frac{t}{2}-\frac{1}{4} \sin (2 t)\text{/.}\, t\to 0\right)=\pi \]